OpenStudy (anonymous):

I need some help with this problem, PLEASE!! 5x-3y=11 and I need to solve for y. I did the math and came up with 11-5x/-3=y but the answer key says the answer is 5x-11/3 =y How is it a positive if you divide by -3?

OpenStudy (phi):

you should get \[ -3y= -5x+11 \\ y = \frac{-5x+11}{-3} \] notice you divide the whole right side by -3 you might remember when you add or subtract fractions, they must have a common denominator example: \[ \frac{2}{3} - \frac{1}{3}= \frac{2-1}{3} \] you can do it backwards: \[ \frac{2-1}{3} =\frac{2}{3} - \frac{1}{3}\] use that idea to write \[ y = \frac{-5x+11}{-3} \\ y = \frac{-5x}{-3} + \frac{11}{-3} \\ y = \frac{5}{3} x - \frac{11}{3} \] or \[ y= \frac{5x - 11}{3} \]

OpenStudy (anonymous):

Oh how simple it seems once you know what you did wrong. Thank you so much...I knew I was just missing something simple. I think I was staring at it too long. :o)